OpenStudy (anonymous):

log 0.000001 simplify if possible

geerky42 (geerky42):

\(10^0 = 1, ~ 10^{-1} = 0.1 ~~\cdots ~~ 10^{-6} = 0.000001\) So \(\log 0.000001 = \log 10^{-6} = -6\log 10 = \boxed{-6}\)

OpenStudy (anonymous):

-6?

geerky42 (geerky42):

Yes.

geerky42 (geerky42):

Is this clear?

OpenStudy (anonymous):

how about log 2.35? approximate the logarithm to four decimal places

geerky42 (geerky42):

you need a calculator to find the answer. Does your calculator have log or ln?

OpenStudy (anonymous):

no unfortunately